Examine linear inequalities and how to formulate their graphical solutions


1
Expert's answer
2021-05-12T06:50:39-0400

In graphical solutions of linear inequalities, solution can be divided into two subsets. 

for example, "2x+y\le 4"


One subset includes all values (x, y) that satisfy the equation 2x+y=4

and the other subset includes all the values (x, y) that satisfy 2x+y<4


First we plot 2x+y=4 on the graph and after that we see the region


By putting (0,0) in the inequality, and see of the conditions true or false-


if condition true- Solution region lies towards the origin.

condition false- solution region lies away from the origin.